OpenStudy (anonymous):

HELP! I don't get this: Solve the system by elimination. -2x + 2y +3z = 0 -2x - y +z = -3 2x + 3y + 3z = 5

OpenStudy (lgbasallote):

look at this: -2x + 2y + 3z = 0 2x + 3y + 3z = 5 you can cancel the x out if you add them right?

OpenStudy (anonymous):

Right. I got that so far.

OpenStudy (lgbasallote):

so what you'll get is 5y + 6z = 5 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

after you get 5y+6z=5, you find 2 other equations from the ORIGINAL 3 and simplify it down so you get another y + z = ___ .

OpenStudy (anonymous):

Thats where I got all mixed up

OpenStudy (anonymous):

so now you pick 2 equations from your original 3 like -2x + 2y +3z = 0 -2x - y +z = -3 and find another y+z=___ to go along with your 5y+6z=5

OpenStudy (anonymous):

but they wont cancel eachother out.

OpenStudy (anonymous):

you can mutliply one of the equations with -1 to switch up the signs

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

I think you get 3y+2z=3? then you just take your 5y+6z=5 and solve for 1 of those variables. then you can practically plug in.

OpenStudy (anonymous):

ok Im trying that right now

OpenStudy (anonymous):

When I multiply -1 to -2x - y + z = -3, I get 2x +y - z = 3, but you said you got 3y+2z=3?

OpenStudy (anonymous):

-2x + 2y +3z = 0 (-2x - y +z = -3)x-1 -2x + 2y +3z = 0 2x +y -z = 3 3y+ 2z = 3

OpenStudy (anonymous):

ohh ok. I am so bad at this

OpenStudy (anonymous):

haha i was confused too when i started this.

OpenStudy (anonymous):

how are you doing on it?

OpenStudy (anonymous):

sorry had to step away. still working on it

OpenStudy (anonymous):

for y I got y=1, if I did it right

OpenStudy (anonymous):

yep i got y=1. now just substitute back into your equation, preferably the y+z, and find z.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so far I have 2 + 4z = 2?

OpenStudy (anonymous):

solve for z

OpenStudy (anonymous):

4z = 0? so z = 0?

OpenStudy (anonymous):

yup! :)

OpenStudy (anonymous):

so you have: z=0, y=1, x=_____

OpenStudy (anonymous):

yay! ok so now just plug y and z back in to x?

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

-2x + 3 = 0 but then what? I subtracted 3 from each side but that didnt work

OpenStudy (anonymous):

which equation did you use?

OpenStudy (anonymous):

oh wait. I just used 2x + 3 + 0 = 5 and got x = 1 is that right?

OpenStudy (anonymous):

pat yourself on the back :) you can go back and check by plugging in all 3 variables in all 3 equations if you want.

OpenStudy (anonymous):

x=1 y=1 z=0

OpenStudy (anonymous):

YAY!!! Thank you so much for your help. I think I finally get the hang of it.